OpenStudy (anonymous):

A rhombus is inscribed in a rectangle that is w meters wide with a perimeter of 24 m. Each vertex of the rhombus is a midpoint of the side of the rectangle. Express the area of the rhombus as a function of the rectangle's width.

OpenStudy (anonymous):



OpenStudy (calculusfunctions):

I would encourage you to have a visual (draw a diagram) to begin.

OpenStudy (calculusfunctions):

OH GOOD!

OpenStudy (calculusfunctions):

Let L represent the rectangles length.

OpenStudy (calculusfunctions):

Now w and L are the width and length respectively. Do you notice there are four congruent right triangles?

OpenStudy (calculusfunctions):

The vertices of the rhombus are located at the midpoints along the width and the length of the rectangle, corect?

OpenStudy (calculusfunctions):

So first, what is the equation for the perimeter of the rectangle, in terms of w and L?

OpenStudy (calculusfunctions):

As a teacher, I make time to come here and teach, and thus am NOT going to just give you the answer so a little participation from you would be greatly appreciated.

OpenStudy (calculusfunctions):

NO?? Well then good luck!

OpenStudy (calculusfunctions):

I'll return when I see you're ready!

OpenStudy (anonymous):

Give me a second.

OpenStudy (anonymous):

@calculusfunctions P=2w+2L

OpenStudy (anonymous):

I do not have the best the connection on here so I may be slow to respond..

OpenStudy (calculusfunctions):

Correct! However since the perimeter is 24 m, then 24 = 2L + 2w OR L + w = 12. Now concentrate on one the four congruent right triangles, and tell me the expression for each side of the rhombus in terms of L and w.

OpenStudy (anonymous):

I am a bit lost.

OpenStudy (calculusfunctions):

Actually, never mind. The area of a rhombus is\[A =\frac{ d _{1}d _{2} }{ 2 }\]where the two d's are the diagonals of the rhombus. Now use this hint to answer the question.

OpenStudy (anonymous):

I am sorry, but I am completely lost on what I am supposed to do next.

OpenStudy (calculusfunctions):

What are the lengths of the diagonals of the rhombus in terms of L and w?

OpenStudy (calculusfunctions):

Look at the diagram Carefully. What the lengths of the diagonals af the rhombus? Do you know what a diagonal is?

OpenStudy (anonymous):

Earlier with the triangle thing did you mean L=w-12and W=L-12

OpenStudy (calculusfunctions):

But then I said "never mind". Don't worry about the triangles.

OpenStudy (anonymous):

And yes I think I do know which the diagonals are

OpenStudy (calculusfunctions):

So then you would see that the diagonals of the rhombus are parallel to the the length and width of the rectangle. Hence the diagonals are L and w. Understand?

OpenStudy (anonymous):

Yes I understand that.

OpenStudy (calculusfunctions):

Thus the area of the rhombus is\[A =\frac{ Lw }{ 2 }\]However you were asked to express the area in terms of w. Thus from the simplified equation (L + w = 12) for the perimeter of the rectangle, from earlier, solve for L and substitute for L into the above area equation.

OpenStudy (calculusfunctions):

Tell me what you get.

OpenStudy (anonymous):

So when you solve for L you get L=12-w and when you substitute it you get A=w(12-w)/2

OpenStudy (calculusfunctions):

Yes! Now simplify.

OpenStudy (anonymous):

I wait isn't \[\frac{ w(w+12) }{ 2 }\] already simplified? or would it be\[\frac{ 1 }{ 2 }w(w+12)\]

OpenStudy (calculusfunctions):

NO! You mean\[A =\frac{ 1 }{ 2 }w(12-w)\]Now simplify.

OpenStudy (calculusfunctions):

Use the distributive property to simplify.

OpenStudy (anonymous):

\[A=6w-\frac{ 1 }{ 2 }w^2\]

OpenStudy (calculusfunctions):

Good! That's correct!

OpenStudy (anonymous):

Thank you for helping!

OpenStudy (calculusfunctions):

Welcome! Anything else?

OpenStudy (anonymous):

I think I'm good for now.